OpenStudy (anonymous):

PLEASE HELPPPP! Consider the function f(x) = 4x+2x^-1. For this function there are four important intervals: (-\infty, A], [A,B),(B,C), and [C,\infty) where A, and C are the critical numbers and the function is not defined at B. Find A,B,C Tell whether the interval is increasing or decreasing

OpenStudy (anonymous):

sorry your screwed I have no idea.

OpenStudy (anonymous):

good thing there are other people who use this site

OpenStudy (anonymous):

@SolomonZelman can you please help me with the question?

OpenStudy (solomonzelman):

Idk how to do this prob

OpenStudy (anonymous):

\[f(x)=4x+2x^{-1}~~\Rightarrow~~f'(x)=4-2x^{-2}=4-\frac{2}{x^2}\] Critical points (note first that \(x=0\) is a critical point since \(f'(x)\) is undefined here): \[\begin{align*} 4-\frac{2}{x^2}&=0\\ 4&=\frac{2}{x^2}\\ 2x^2-1&=0\\\\ x^2-\frac{1}{2}&=0\\\\ x&=\pm\frac{1}{\sqrt2} \end{align*}\] So the intervals would be \[\left(-\infty,-\frac{1}{\sqrt2}\right)\\ \left(-\frac{1}{\sqrt2},0\right)\\ \left(0,\frac{1}{\sqrt2}\right)\\ \left(\frac{1}{\sqrt2},\infty\right)\]

OpenStudy (solomonzelman):

DDDDDang ... the solutions looks wise :)